OpenStudy (anonymous):

An outside circular ring has a circumference of 90 cm. What is the circumference of an inner ring which is 8 cm shorter in radius? Both circles have the same center. i came to 39.8but im barley learning this so just making sure if its wrong don't need to explain just say its wrong thx

OpenStudy (anonymous):



OpenStudy (anonymous):

\[{\large{perimeter = 2\times{pie}\times{r}}}\]

OpenStudy (anonymous):

@sheg means perimeter of circle. :D And one latex tip to write pi. Write as \pi

OpenStudy (anonymous):

@shivam_bhalla thanks a ton buddy

OpenStudy (anonymous):

oqei thx

OpenStudy (anonymous):

Your approach towards the problem must be to find the radius of outer circle using the formula given by @sheg and then......... @sheg , welcome :)